[Brahms-l] Shift report 20040223 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Mon Feb 23 2004 - 07:55:17 EST
------------------------------------------------------------
Date: Mon Feb 23 07:55:16 2004, Shift: 00:00-08:00
Supervisor: DO, with: 
Log book#15 pages 65 - 65

Shift Summary:
------------------------------------------------------------
1. store from previous shift until 1.30am
   2nd store started at 2am until 6.30am
   3rd store started at ~ 7am, still going -> 7h

2. Data:
   MRS 40 B1050: 583k tr3 (runs 10525 -> 10538)
   FS  4  A1/2 : 539k tr2
   run 10549 still going

   pedestals:
   non TPC: 10526 10539
   TPC ped: 10529 10540

3. A few problems that could be fixed:
   At shift start, noticed badtrip was off. A close
   inspection of TPC HVs showed that TPM1 anode and 
   drift HV were off. Mistake from previous shifters.
   Brought TPM1 on, traced back bad runs from previous 
   shift (10501, 10523, 10524), changed MRS quality
   accordingly.
   Lesson to learn: look at online summary plots! Note that
   TPC plots are not included in the autoprint, it could
   have eased the search for bad runs.

   Second problem between stores 1 and 2: 
   io and run servers screwed up. Then trigger counts 
   not incremented properly. Reset all servers (io, run,
   monitorbeam, monitorscaler). After 2-3 times, it 
   worked ok.

   Same thing between stores 2 & 3. I found out that it is
   important to do things in the rigth order, namely:
   stop io server, stop run server, start run server,
   start io server. Otherwise, things remain screwed,
   supermon cannot connect to event builder, etc.

   Note again that event rate shows oscillations, sometimes 
   the DAQ pauses for 3-4 seconds.

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Feb 23 07:56:37 2004

This archive was generated by hypermail 2.1.8 : Mon Feb 23 2004 - 07:56:57 EST